# Natural Metric-Affine Inflation

### v1

@@ Abstract
We consider here natural inflation in the low energy (two-derivative) metric-affine theory containing only the minimal degrees of freedom in the inflationary sector, i.e. the massless graviton and the pseudo-Nambu-Goldstone boson (PNGB). This theory contains the Ricci-like and parity-odd Holst invariants together with non-minimal couplings between the PNGB and the above-mentioned invariants. The Palatini and Einstein-Cartan realizations of natural inflation are particular cases of our construction. Explicit models of this type are shown to admit an explicit UV completion in a QCD-like theory with a Planckian confining scale. Moreover, for these models, we find regions of the parameter space where the inflationary predictions agree with the most recent observations at the $2\sigma$ level. We find that in order to enter the $1\sigma$ region it is necessary (and sufficient) to have a finite value of the Barbero-Immirzi parameter and a sizable non-minimal coupling between the inflaton and the Holst invariant (with sign opposite to the Barbero-Immirzi parameter). Indeed, in this case the potential of the canonically normalized inflaton develops a plateau as shown analytically.
@@ 2 The action and the dynamical degrees of freedom
where here $\phi$ is the NBGB field,
@@
where $R$ is the Ricci scalar.
@@
In the case of natural inflation formulated directly in the metric formalism, namely starting from the action
@@
This $k(\phi)$ differs from the one in (2.14) because of the extra positive term $3M_{P}^{2}F^{\prime 2}(\phi)/(2F^{2}(\phi))$.
@@ 3 An analytical study of the inflationary potential
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ 4 Inflationary predictions
$\epsilon(\chi)$ $=$ $\frac{M_{P}^{2}}{2}\left(\frac{U^{\prime}(\chi)}{U(\chi)}\right)^{2}\,,$ (4.1)
$\eta(\chi)$ $=$ $M_{P}^{2}\frac{U^{\prime\prime}(\chi)}{U(\chi)}\,.$ (4.2)
@@
where the field value at the end of inflation is given by $\epsilon(\chi_{\textrm{end}})=1$, while the field value $\chi_{N}$ at the time a given scale left the horizon is given by the corresponding $N_{e}$. The spectral index $n_{\textrm{s}}$ and the tensor-to-scalar ratio $r$ are:
$n_{\textrm{s}}$ $=$ $1+2\eta(\chi_{N})-6\epsilon(\chi_{N})\,,$ (4.4)
$r$ $=$ $16\epsilon(\chi_{N})\,.$ (4.5)
@@
$A_{\textrm{s}}=\frac{1}{24\pi^{2}M_{P}^{4}}\frac{U(\chi_{N})}{\epsilon(\chi_{N})}\,,$ (4.6)
@@ 4.1.1 $\xi<0$ and $\beta(\phi)=0$
In Fig. 4 we plot the corresponding results for $r$ and $n_{s}$ versus the parameters $\xi$ and $\delta_{f}$ when $N_{e}=60$ and $\delta_{\Lambda}$ is fixed so that the constraint (4.7) is satisfied.
We can see that with $\xi$ increasing in absolute value $r$ ($n_{s}$) increases (decreases) leading the prediction even more away from the allowed region. This was somehow expected because a negative $\xi$ increases the height of maximum of $U$ in Eq. (3.8) and this usually comes with an increase in $r$. Therefore, we conclude that the configuration with $\xi<0$ and $\beta(\phi)=0$ is excluded by data.
@@ 4.1.2 $\xi>0$ and $\beta(\phi)=0$
In Fig. 5 we plot the corresponding results for $r$ and $n_{s}$ as functions of the parameters $\xi$ and $\delta_{f}$ when $N_{e}=60$ and $\delta_{\Lambda}$ is fixed so that the constraint (4.7) is satisfied.
Regardless of $\delta_{f}$, the predictions share some common properties. First, as expected, $r$ decreases by increasing $\xi$, while $n_{s}$ first reaches a maximum (whose numerical value is dependent on the actual value of $\delta_{f}$) and then decreases. A similar behaviour is also shown by $\delta_{\Lambda}$ as a function of $\xi$. Moreover, $\delta_{\Lambda}$ also increases by increasing $\delta_{f}$. At selected $\delta_{f}$ values, by increasing $\xi$ the predictions manage to reach the $2\sigma$ allowed region. In particular for $\delta_{f}=6$, they touch the border of the $1\sigma$ region. This happens for $\xi$ values around 1 or smaller. For $\delta_{f}\gtrsim 12$, no predictions enter the allowed region.
The results can be better understood by giving a look at the Einstein frame potential $U(\chi)$, which is plotted in Fig. 6 for selected benchmark points. For clarity the lines have been truncated at half period i.e. $\chi=\chi(\phi_{2})$. First of all we can see that the potential is lowered and flattened as an effect of the non-minimal coupling. This usually comes with a lowering of $r$, which is confirmed by the predictions in Fig. 5. We also see that by increasing $\delta_{f}$ (at fixed $\xi$), $\chi_{N}$ moves further and further away from the maximum of the potential towards steeper regions, implying an increase on $r$, which is confirmed by Fig. 5b.
Overall we can conclude that the presence of the non-minimal coupling $\xi$ is beneficial to restore compatibility with data of natural inflation. This is also the case in the metric theory of natural inflation in the presence of the non-minimal coupling [11].
@@ 4.2.1 $\beta_{0}\geq 0$
In Fig. 7 we plot the corresponding results for $r$ and $n_{s}$ versus the parameters $\tilde{\xi}$ and $\delta_{f}$ when $\beta_{0}\geq 0$, $N_{e}=60$ and $\delta_{\Lambda}$ is fixed so that the constraint (4.7) is satisfied.
We can see that the $\beta_{0}>0$ and $\beta_{0}=0$ results overlap with each other. Therefore, the effect of having $\beta_{0}>0$ is just to shift the results at higher $\tilde{\xi}$ values. Unfortunately, with $\tilde{\xi}$ increasing in both cases $r$ ($n_{s}$) increases (decreases) leading the prediction even more away from the allowed region. Because of this, we truncated the study at $\tilde{\xi}=10000$ and concluded that the configuration with $\xi=0$ and $\beta_{0}\geq 0$ is excluded by data.
@@ 4.2.2 $\beta_{0}<0$
Figure 9: $U(\chi)$ (a) and $U(\chi-\chi_{\text{flex}})$ (b) with $\xi=0$, $\beta_{0}=-6M_{P}^{2}$ and $n_{s}\simeq 0.97$ (continuous) and $U(\chi)$ with $\xi=\beta_{0}=\tilde{\xi}=0$ (dashed) for $\delta_{f}$ ranging from 4 (purple) to 10 (red) with steps of 1, displayed in rainbow colors. In the continuous lines $\tilde{\xi}$ varies with $f$ such that $n_{s}\simeq 0.97$. The bullets represents the corresponding points at $\chi_{N}$ with $N_{e}=60$.
In Fig. 8 we plot the corresponding results for $r$ and $n_{s}$ versus the parameters $\tilde{\xi}$ and $\delta_{f}$ when $\beta_{0}=-6M_{P}^{2}$, $N_{e}=60$ and $\delta_{\Lambda}$ is fixed so that the constraint (4.7) is satisfied.
We can see that by increasing $\tilde{\xi}$, $r$ decreases in most of the chosen configuration but $\delta_{f}=4,5$ where, instead, $r$ vs. $\tilde{\xi}$ exhibits first an increasing behaviour and then a decreasing one like in all the other $\delta_{f}$’s. On the other hand, the behaviour of $n_{s}$ vs. $\tilde{\xi}$ is more oscillating (showing first a short increasing phase, then a decreasing one and finally a new increasing phase) for all the chosen configurations but for $\delta_{f}=4$ where $n_{s}$ appears to just increase with $\tilde{\xi}$. At big enough $\tilde{\xi}$, the results in the region $r\lesssim 0.015$ and $n_{s}\gtrsim 0.958$ overlap, but they are still dependent on $f$. This is visible in Fig. 8(c). This means that by changing $f$, we can find a $\tilde{\xi}$ so that the results instead do not change.
To get a better understanding, we plot for selected benchmark points $U(\chi)$ centered at the origin and centered around the corresponding inflection point respectively in Fig. 9(a) and in Fig. 9(b). For clarity the lines have been truncated at half period i.e. $\chi=\chi(\phi_{2})$. For the selected points, $\chi_{N}$ is close to the inflection point of the potential. Moreover, it is hard to distinguish the different potentials around (and after) the inflection points. The equation for such a points is
@@ 4.3 $\xi>0$ and $\tilde{\xi}>0$
In Fig. 10 we plot the corresponding results for $r$ and $n_{s}$ versus the parameters $\tilde{\xi}$ and $\delta_{f}$ when $\xi=\frac{1}{3}$, $\beta_{0}=-2M_{P}^{2}$, $N_{e}=60$ and $\delta_{\Lambda}$ is fixed so that the constraint (4.7) is satisfied.
The results are clearly a combination of the behaviour pattern shown in Figs. 5 and 8. As expected compatibility with data is achieved quite easily.

### v2

@@ Abstract
We consider here natural inflation in the low energy (two-derivative) metric-affine theory containing only the minimal degrees of freedom in the inflationary sector, i.e. the massless graviton and the pseudo-Nambu-Goldstone boson (PNGB). This theory contains the Ricci-like and parity-odd Holst invariants together with non-minimal couplings between the PNGB and the above-mentioned invariants. The Palatini and Einstein-Cartan realizations of natural inflation are particular cases of our construction. Explicit models of this type featuring non-minimal couplings are shown to emerge from the microscopic dynamics of a QCD-like theory with an either sub-Planckian or trans-Planckian confining scale and that is renormalizable on Minkowski spacetime. Moreover, for these models, we find regions of the parameter space where the inflationary predictions agree with the most recent observations at the $2\sigma$ level. We find that in order to enter the $1\sigma$ region it is necessary (and sufficient) to have a finite value of the Barbero-Immirzi parameter and a sizable non-minimal coupling between the inflaton and the Holst invariant (with sign opposite to the Barbero-Immirzi parameter). Indeed, in this case the potential of the canonically normalized inflaton develops a plateau as shown analytically.
@@ 2 The action and the dynamical degrees of freedom
where here $\phi$ is the PNGB field,
@@
where $R$ is the Ricci scalar (remember that $R=\cal R$ when ${\cal A}_{\mu\leavevmode\nobreak\ \sigma}^{\leavevmode\nobreak\ \,\rho}=\Gamma_{\mu\leavevmode\nobreak\ \sigma}^{\leavevmode\nobreak\ \,\rho}$).
@@
In the case of natural inflation formulated directly in the metric formalism (as opposed to the metric-affine formalism), namely starting from the action
@@
This $k(\phi)$ differs from the one in (2.14) because of the extra positive term $3M_{P}^{2}F^{\prime 2}(\phi)/(2F^{2}(\phi))$. Therefore, we can see that the two theories differ even if one sets $\beta=0$ in the metric-affine case.
@@ 3 An analytical study of the inflationary potential
Before ending this section let us observe that in Fig. 2 we have chosen trans-Planckian values of $f$. As we will see in Sec. 4, among others, it is also for such values that we reach the agreement with CMB observations. Nevertheless, it is important to note that this setup do not necessarily invalidate the effective-field-theory treatment of gravity, which we are relying on. Indeed, the condition that allows us to do so is that the energy density should be much less than the Planck energy density [35]
$|U(\chi)|\ll M_{P}^{4},$ (3.14)
for all relevant values of $\chi$. As we will show in Sec. 4, this condition is amply satisfied. It should be kept in mind, however, than in some UV completions of all interactions such as string theory [36] having a trans-Planckian $f$ can spoil the validity of natural inflation, at least in its simplest incarnation [37, 38] (see [39] for a review). In some others, such as asymptotically free/safe theories [40] featuring higher-derivative terms (see e.g. [41] and references therein) a trans-Planckian $f$ can be consistent [8]. We find this aspect interesting as observing the predictions of natural inflation may give us information on the theory that UV completes Einstein gravity. For this reason, in Sec. 4 we will consider both trans-Planckian and sub-Planckian values of $f$ and, as we will show, agreement with data can be achieved in both cases.
@@ 4 Inflationary predictions
$\epsilon_{U}(\chi)$ $=$ $\frac{M_{P}^{2}}{2}\left(\frac{U^{\prime}(\chi)}{U(\chi)}\right)^{2}\,,$ (4.1)
$\eta_{U}(\chi)$ $=$ $M_{P}^{2}\frac{U^{\prime\prime}(\chi)}{U(\chi)}\,,$ (4.2)
$\xi^{2}_{U}(\chi)$ $=$ $M_{P}^{4}\frac{U^{\prime}(\chi)U^{\prime\prime\prime}(\chi)}{U(\chi)^{2}}\,.$ (4.3)
@@
where the field value at the end of inflation is given by $\epsilon(\chi_{\textrm{end}})=1$, while the field value $\chi_{N}$ at the time a given scale left the horizon is given by the corresponding $N_{e}$. The tensor-to-scalar ratio $r$, the spectral index $n_{\textrm{s}}$ and its running $\alpha_{\text{s}}\equiv\text{d}n_{\text{s}}/\text{d}\ln k$ are:
$r$ $=$ $16\epsilon_{U}(\chi_{N})\,,$ (4.5)
$n_{\textrm{s}}$ $=$ $1+2\eta_{U}(\chi_{N})-6\epsilon_{U}(\chi_{N})\,,$ (4.6)
$\alpha_{\textrm{s}}$ $=$ $16\epsilon_{U}(\chi_{N})\eta_{U}(\chi_{N})-24\epsilon_{U}^{2}(\chi_{N})-2\xi^{2}_{U}(\chi_{N})\,.$ (4.7)
@@
$A_{\textrm{s}}=\frac{1}{24\pi^{2}M_{P}^{4}}\frac{U(\chi_{N})}{\epsilon_{U}(\chi_{N})}\,,$ (4.8)
@@ 4.1.1 $\xi<0$ and $\beta(\phi)=0$
In Fig. 4 we plot the corresponding results for $r$ and $n_{\text{s}}$ versus the parameters $\xi$ and $\delta_{f}$ when $N_{e}=60$ and $\delta_{\Lambda}$ is fixed so that the constraint (4.9) is satisfied.
We can see that with $\xi$ increasing in absolute value $r$ ($n_{\text{s}}$) increases (decreases) leading the prediction even more away from the allowed region. This was somehow expected because a negative $\xi$ increases the height of maximum of $U$ in Eq. (3.8) and this usually comes with an increase in $r$. Therefore, we conclude that the configuration with $\xi<0$ and $\beta(\phi)=0$ is excluded by data and the computation of the predictions for $\alpha_{\text{s}}$ is not needed.
@@ 4.1.2 $\xi>0$ and $\beta(\phi)=0$
In Fig. 5 we plot the corresponding results for $r$ and $n_{\text{s}}$ as functions of the parameters $\xi$ and $\delta_{f}$ when $N_{e}=60$ and $\delta_{\Lambda}$ is fixed so that the constraint (4.9) is satisfied.
Regardless of $\delta_{f}$, the predictions share some common properties. First, as expected, $r$ decreases by increasing $\xi$, while $n_{\text{s}}$ first reaches a maximum (whose numerical value is dependent on the actual value of $\delta_{f}$) and then decreases. A similar behaviour is also shown by $\delta_{\Lambda}$ as a function of $\xi$. Moreover, $\delta_{\Lambda}$ also increases by increasing $\delta_{f}$. At selected $\delta_{f}$ values, by increasing $\xi$ the predictions manage to reach the $2\sigma$ allowed region. In particular for $\delta_{f}=6$, they touch the border of the $1\sigma$ region. This happens for $\xi$ values around 1 or smaller. For $\delta_{f}\gtrsim 12$, no predictions enter the allowed region.
The results can be better understood by giving a look at the Einstein frame potential $U(\chi)$, which is plotted in Fig. 7 for selected benchmark points. For clarity the lines have been truncated at half period i.e. $\chi=\chi(\phi_{2})$. First of all we can see that the potential is lowered and flattened as an effect of the non-minimal coupling. As anticipated in Sec. 3, the condition in (3.14) to rely on an effective-field-theory treatment of gravity is amply satisfied (see Fig. 7). The main reason why this happens is the smallness of $\Lambda$ in Planck units, $\delta_{\Lambda}\ll 1$, (shown in Fig. 5d); in the construction of Appendix A such smallness is natural because it is generated by the smallness of quark-like masses (the only source of a chiral symmetry breaking) compared to a QCD-like confining scale. The flattening of the potential usually comes with a lowering of $r$, which is confirmed by the predictions in Fig. 5. We also see that by increasing $\delta_{f}$ (at fixed $\xi$), $\chi_{N}$ moves further and further away from the maximum of the potential towards steeper regions, implying an increase on $r$, which is confirmed by Fig. 5b.
[…]
Overall we can conclude that the presence of the non-minimal coupling $\xi$ is beneficial to restore compatibility with data of natural inflation. This is also the case in the metric theory of natural inflation in the presence of the non-minimal coupling [11], where $k(\phi)$ is given by (2.16) rather than (2.14).
@@ 4.2.1 $\beta_{0}\geq 0$
In Fig. 8 we plot the corresponding results for $r$ and $n_{\text{s}}$ versus the parameters $\tilde{\xi}$ and $\delta_{f}$ when $\beta_{0}\geq 0$, $N_{e}=60$ and $\delta_{\Lambda}$ is fixed so that the constraint (4.9) is satisfied.
We can see that the $\beta_{0}>0$ and $\beta_{0}=0$ results overlap with each other. Therefore, the effect of having $\beta_{0}>0$ is just to shift the results at higher $\tilde{\xi}$ values. Unfortunately, with $\tilde{\xi}$ increasing in both cases $r$ ($n_{\text{s}}$) increases (decreases) leading the prediction even more away from the allowed region. Because of this, we truncated the study at $\tilde{\xi}=10000$ and concluded that the configuration with $\xi=0$ and $\beta_{0}\geq 0$ is excluded by data and the computation of the predictions for $\alpha_{\text{s}}$ is not needed.
@@ 4.2.2 $\beta_{0}<0$
Figure 10: $\alpha_{\text{s}}$ vs. $n_{\text{s}}$ (a), $\alpha_{\text{s}}$ vs. $\tilde{\xi}$ (b) for $N_{e}=60$ and $r\lesssim 0.0383$ with $\xi=0$, $\beta_{0}=-6M_{P}^{2}$ for the same $\delta_{f}$ values and color code as in Fig. 9. The pink and gray areas are the same as in Fig. 6.
In Fig. 9 we plot the corresponding results for $r$ and $n_{\text{s}}$ versus the parameters $\tilde{\xi}$ and $\delta_{f}$ when $\beta_{0}=-6M_{P}^{2}$, $N_{e}=60$ and $\delta_{\Lambda}$ is fixed so that the constraint (4.9) is satisfied.
Figure 11: $U(\chi)$ (a) and $U(\chi-\chi_{\text{flex}})$ (b) with $\xi=0$, $\beta_{0}=-6M_{P}^{2}$ and $n_{\text{s}}\simeq 0.97$ (continuous) and $U(\chi)$ with $\xi=\beta_{0}=\tilde{\xi}=0$ (dashed) for $\delta_{f}$ ranging from 4 (purple) to 10 (red) with steps of 1, displayed in rainbow colors. In the continuous lines $\tilde{\xi}$ varies with $f$ such that $n_{\text{s}}\simeq 0.97$. The bullets represents the corresponding points at $\chi_{N}$ with $N_{e}=60$.
We can see that by increasing $\tilde{\xi}$, $r$ decreases in most of the chosen configuration but $\delta_{f}=4,5$ where, instead, $r$ vs. $\tilde{\xi}$ exhibits first an increasing behaviour and then a decreasing one like in all the other $\delta_{f}$’s. On the other hand, the behaviour of $n_{\text{s}}$ vs. $\tilde{\xi}$ is more oscillating (showing first a short increasing phase, then a decreasing one and finally a new increasing phase) for all the chosen configurations but for $\delta_{f}=4$ where $n_{\text{s}}$ appears to just increase with $\tilde{\xi}$. At big enough $\tilde{\xi}$, the results in the region $r\lesssim 0.015$ and $n_{\text{s}}\gtrsim 0.958$ overlap, but they are still dependent on $f$. This is visible in Fig. 9(c). This means that by changing $f$, we can find a $\tilde{\xi}$ so that the results instead do not change.
Once again, it is worth studying the predictions for $\alpha_{\text{s}}$. In Fig. 10 we plot the corresponding results as functions of the parameters $\tilde{\xi}$ and $\delta_{f}$. The same cut in $r$ as in Fig. 6 has been implemented. We can see that the constraint on $\alpha_{\text{s}}$ reduces the allowed parameter space. Again all the predicted values for $\alpha_{\text{s}}$ are negative but now larger in absolute value than those of the $\beta=0$ case in Fig. 6.
To get a better understanding of the large $\tilde{\xi}$ behaviour, we plot for selected benchmark points $U(\chi)$ centered at the origin and centered around the corresponding inflection point respectively in Fig. 11(a) and in Fig. 11(b). For clarity the lines have been truncated at half period i.e. $\chi=\chi(\phi_{2})$. For the selected points, $\chi_{N}$ is close to the inflection point of the potential. Moreover, it is hard to distinguish the different potentials around (and after) the inflection points. The equation for such a points is
@@
As anticipated in Sec. 3, once again we note that the condition in (3.14) to rely on an effective-field-theory treatment of gravity is amply satisfied, as shown by Fig. 11. Also here the main reason why this happens is the smallness of $\delta_{\Lambda}$, (shown here in Fig. 9d), which is natural in the construction of Appendix A.
@@
Figure 13: $\alpha_{\text{s}}$ vs. $n_{\text{s}}$ (a), $\alpha_{\text{s}}$ vs. $\tilde{\xi}$ (b) for $N_{e}=60$ and $r\lesssim 0.0383$ with $\xi=\frac{1}{3}$, $\beta_{0}=-2M_{P}^{2}$ for the same $\delta_{f}$ values and color code of Fig. 12. The pink and gray areas are the same as in Fig. 6.
@@ 4.3 $\xi>0$ and $\tilde{\xi}>0$
Figure 14: $r$ vs. $n_{\text{s}}$ (a), $r$ vs. $\tilde{\xi}$ (b), $\tilde{\xi}$ vs. $n_{\text{s}}$ (c), $\delta_{\Lambda}$ vs. $\tilde{\xi}$ (d) for $N_{e}=60$ with $\beta_{0}=-10M_{P}^{2}$ for $\delta_{f}=10^{-2}$ (purple), $\delta_{f}=10^{-1}$ (yellow) and $\delta_{f}=1$ (red) when $\xi=\frac{1}{3}$ (continuous) or $\xi=0$ (dashed). The gray color codes are the same as in Fig. 4.
Figure 15: $\alpha_{\text{s}}$ vs. $n_{\text{s}}$ (a), $\alpha_{\text{s}}$ vs. $\tilde{\xi}$ (b) for $N_{e}=60$ with $\beta_{0}=-10M_{P}^{2}$ for the same values and color codes of $\delta_{f}$ and $\xi$ as in Fig. 14. The pink and gray areas are the same as in Fig. 6.
In Figs. 12 and  13 we plot the corresponding results for $r$, $n_{\text{s}}$ and $\alpha_{\text{s}}$ versus the parameters $\tilde{\xi}$ and $\delta_{f}$ when $\xi=\frac{1}{3}$, $\beta_{0}=-2M_{P}^{2}$, $N_{e}=60$ and $\delta_{\Lambda}$ is fixed so that the constraint (4.9) is satisfied.
The results are clearly a combination of the behaviour pattern shown in subsections 4.1.2 and 4.2.2. The most interesting result is the appearance of regions where a positive $\alpha_{\text{s}}$ is predicted. However, these regions are ruled out because of the constraint on $r$ vs. $n_{\text{s}}$, but for $\delta_{f}=3$ where the positive $\alpha_{\text{s}}$ is excluded by the $\alpha_{\text{s}}$ vs. $n_{\text{s}}$ constraint.
Overall, as expected, compatibility with data is achieved quite easily.
Moreover, this case is interesting also because it allows us to satisfy the experimental constraints even when $f$ is sub-Planckian. The corresponding results are shown in Figs. 14 and 15. For numerical convenience the study considers only values of $\tilde{\xi}$ between approximately 10 and 20. The lines representing to the results corresponding to $\delta_{f}=10^{-2}$ and $\delta_{f}=10^{-1}$ are essentially undistinguishable apart for the $\alpha_{s}$ vs. $n_{s}$ plot. We can see that when $\xi=0$, with a large enough $\tilde{\xi}$, the $r$ vs. $n_{s}$ predictions at very low $\delta_{f}$ are again compatible with data. However, in this case $\alpha_{s}$ is very large and essentially always out from the allowed region for sub-Planckian $f$. This was somehow expected because in standard natural inflation the second slow-roll parameter $\eta$ is very large in absolute value at low $\delta_{f}$ (potentially even breaking the slow-roll approximation), implying a very small $n_{s}$. In order to achieve compatibility with data, $n_{s}$ is varying very fast causing its running $\alpha_{s}$ to be out of the allowed region. The issue can be solved by using $\xi>0$, which restores compatibility with data by moving the large $\alpha_{s}$ region towards smaller $n_{s}$ values.